OpenStudy (anonymous):

trace the curve y^2 = x^2 + y - 6

OpenStudy (anonymous):

Algebraic curve

OpenStudy (anonymous):

\[y^2-y-x^2=-6\] \[(y-1/2)^2-x^2=1/4-6=-23/4\] \[x^2-(y-\frac{1}{2})^2=\frac{23}{4}\] looks like a hyperbolar

OpenStudy (ankit042):

@jonask How about a circle with center at (0,1/2) and radius of sqrt(23/4) (x-a)^2+(y-b)^2 =r^2 standard equation of a circle

OpenStudy (anonymous):

these a minus between the brackets so its a problem (x-a)^2-(y-b)^2 =r^2

OpenStudy (ankit042):

ohh Crap! sorry I didn't see your solution properly :(